# Medium separation scheme effects on the magnetized and cold two-flavor superconducting quark matter

### v1

@@ Abstract
We analyze the impact of the Medium Separation Scheme (MSS) on two-flavor color superconducting (2SC) dense quark matter under the influence of a constant external magnetic field. The effects of the proper treatment of the model divergences are examined through a comparison of different approaches, including the combined implementation of the Magnetic Field Independent Regularization (MFIR) and the MSS, as well as the standard use of smooth form factors. Our findings for the Nambu–Jona-Lasinio model emphasize the critical role of properly separating medium effects from vacuum contributions in the model. The combined MFIR-MSS scheme suppresses spurious unphysical oscillations, often misinterpreted in the literature as de Haas–van Alphen oscillations, and ensures the correct high-density behavior of the diquark condensate. Furthermore, within the MSS framework, the magnetization remains positive across the explored parameter space, in sharp contrast with the behavior obtained in the traditional approach.
@@ I Introduction
The possible existence of deconfined quark matter in the core of neutron stars Annala et al. (2020) has shed new light on the understanding of strongly interacting matter described by Quantum Chromodynamics (QCD) at high densities. The expectation that quark matter becomes color superconducting at very high densities has been thoroughly reviewed in the literature Bailin and Love (1984). Perturbative calculations were later confirmed within a controlled weak-coupling framework in Ref. Son (1999), where the authors derived the parametric form of the superconducting gap. Nevertheless, these perturbative approaches are not applicable to the intermediate density regime, which is relevant for the phenomenology of compact objects. Therefore, alternative approaches must be developed to explore the properties of matter in this regime. From a theoretical perspective, much of our current understanding of the QCD phase structure largely stems from first-principles Monte Carlo simulations on the lattice Laermann and Philipsen (2003); Parotto (2023), although the well-known sign problem limits the applicability of this approach at finite chemical potential (see Nagata (2022) for a recent review). Consequently, our understanding of the behavior of cold and dense matter in the intermediate density regime relies heavily on the development of effective models. Chiral models such as the Nambu–Jona-Lasinio (NJL) Klevansky (1992); Buballa (2005) and the linear sigma model Ayala et al. (2015); Andersen and Nødtvedt (2025) aim to capture the essential features of strongly interacting matter by incorporating key symmetries and physical constraints of QCD, providing valuable insights into the equation of state and phase transitions relevant to compact stars and other dense systems.
@@
Color superconducting phases are particularly relevant in astrophysical contexts such as magnetars, where dense matter is subjected to strong magnetic fields. In these extreme environments, the standard electromagnetic Meissner effect may be absent for certain non-spherical superconducting phases, thereby allowing the magnetic flux to penetrate without significant energy cost Feng et al. (2010). Earlier investigations of magnetized color-superconducting quark matter have demonstrated that strong magnetic fields can significantly influence both pairing dynamics and the phase structure, modifying the behavior of the gap and the equation of state Abhishek and Mishra (2019). Within the two-flavor sector, rotated charges and anisotropic superconductivity have been investigated in Refs. Coppola et al. (2017); Cao and Zhuang (2015), emphasizing the necessity of incorporating electric and color neutrality constraints for more realistic descriptions. Furthermore, intense magnetic fields have been shown to induce oscillatory behavior in the gap parameters Fukushima and Warringa (2008a); Noronha and Shovkovy (2007), emphasizing the non-trivial interplay between magnetic fields and superconductivity. Beyond gap oscillations, more exotic phenomena may arise, such as the formation of inhomogeneous gluon condensates, which mitigate chromomagnetic instabilities in neutral two-flavor superconductors, and the emergence of rotated magnetic fields Ferrer and de la Incera (2006, 2007a); Yuan et al. (2024). These effects are crucial for accurately modeling compact stars, especially magnetars. In the three-flavor case, intense magnetic fields may induce new phases, such as the Magnetic Color-Flavor-Locked (MCFL) and Paramagnetic-Color-Flavor-Locked (PCFL) phases, wherein both the pairing structure and the spectrum of collective excitations are significantly modified Ferrer and de la Incera (2007b); Feng et al. (2011); Paulucci et al. (2011). The alignment of Cooper pair magnetic moments in these phases can lead to the formation of additional condensates, thereby enhancing the pairing energy in the presence of strong fields. From a theoretical perspective, renormalization-group analyses have elucidated the role of long-range magnetic interactions in color superconductivity at asymptotically high densities, leading to a non-BCS scaling of the superconducting gap, $\Delta\sim\mu\,g^{-5}\exp(-c/g)$ Son (1999); Hsu and Schwetz (2000), where $c$ is a constant. These insights are essential for the construction of consistent models of dense QCD matter under extreme magnetic conditions. Moreover, it has been proposed that strong background fields could induce an electromagnetic superconducting state of the vacuum through the condensation of charged vector $\rho$ mesons, suggesting an analogy between QCD and condensed matter systems Chernodub (2010, 2013). However, recent calculations based on extended NJL models Carlomagno et al. (2022a, b), which consistently include Schwinger phases and proper wave functions for $B\neq 0$, have found no evidence of such condensation in the range of magnetic fields considered, thus disfavoring the occurrence of this mechanism.
@@
In this work, we employ the NJL model to study diquark condensation in cold and dense quark matter under an external magnetic field. Given the nonrenormalizable nature of the NJL model, the regularization procedure must be handled with particular care. In this class of effective models, ultraviolet (UV) divergent momentum integrals must be handled, typically by introducing a sharp cutoff $\Lambda$, which is regarded as a model parameter fitted to physical observables. This cutoff defines an energy scale beyond which the model cannot be trusted. Standard procedures often mix vacuum and medium effects, introducing dependencies on chemical potential, temperature, and external fields. Given that the NJL model serves as an effective approximation of QCD, it is reasonable to impose that regularization affects only vacuum-dependent terms. This principle was first applied in studies of color superconductivity Farias et al. (2006) and has since been employed in different contexts to describe the QCD phase diagram. The separation, termed the Medium Separation Scheme (MSS), contrasts with the traditional regularization scheme (TRS), which applies the cutoff to all divergent integrals that include physical quantities that depend on medium effects. One of the main motivations for this study is the evidence that the diquark condensate $\Delta$ increases with baryon chemical potential $\mu_{B}$. While lattice QCD simulations for $N_{c}=3$ are hindered by the sign problem Karsch (2002), studies for $N_{c}=2$ Kogut et al. (2001); Braguta et al. (2016), Chiral Perturbation Theory (ChPT) Kogut et al. (2000) and the renormalization-group consistent treatment applied to the NJL model Gholami et al. (2025) show this increasing behavior of $\Delta$ with $\mu_{B}$. Using TRS, the diquark condensate starts to decrease at a value of chemical potential smaller than the energy scale $\Lambda$, and eventually vanishes. In contrast, employing MSS, a consistently increasing condensate is predicted, in agreement with other approaches.
@@
Over the recent past, significant attention has been devoted to the generation of magnetic fields in noncentral heavy-ion collisions, where the field strength can reach up to $10^{20}$ G, and which can significantly influence the QCD phases Bzdak and Skokov (2013); McLerran and Skokov (2014). The study of QCD under strong magnetic fields has expanded across different contexts, including magnetars, neutron star mergers, and early universe physics. A recent review Adhikari and others (2024) (and references therein) presents a thorough discussion of the different contexts where magnetic fields can influence the phase structure and transitions of QCD matter. In this context, accurately modeling the behavior of quark matter under strong magnetic fields requires a careful choice of regularization schemes to avoid unphysical artifacts and ensure reliable predictions, especially when working with nonrenormalizable models. The choice of an appropriate regularization scheme is essential for accurately describing magnetized quark matter, as highlighted in previous studies Avancini et al. (2019b); Allen et al. (2015); Duarte et al. (2016, 2017). As an example, one may cite a recent work Endrődi and Markó (2019), where the authors fixed the running coupling of the two-flavor Polyakov extended NJL model (PNJL) by matching its predictions with LQCD results for the baryon spectrum in a strong magnetic background. This approach, which preserves the expected decrease of the pseudocritical temperature, $T_{pc}$, with the increase of the magnetic field for both chiral and deconfinement transitions, relies on Schwinger’s proper time formalism Schwinger (1951), where all integrals are regularized without separating vacuum and thermomagnetic contributions. However, while this method ensures consistency with LQCD predictions at moderate temperatures, it can lead to inconsistencies at higher temperatures or finite baryon chemical potentials, such as unphysical oscillations in magnetization and imaginary meson masses Costa et al. (2010a); Fayazbakhsh et al. (2012); Ruivo et al. (2010); Costa et al. (2010b). Different regularization schemes have been proposed to address these issues, with the most prominent being the Magnetic Field Independent Regularization (MFIR), proposed in Ebert et al. (1999); Ebert and Klimenko (2003) and applied in Menezes et al. (2009), which isolates vacuum divergences from medium contributions and removes unphysical oscillations Ebert and Klimenko (2003); Ebert et al. (1999). Additionally, variations in handling the thermal contribution, such as the standard proper time and the thermomagnetic regulated proper time schemes Endrődi and Markó (2019), affect the convergence to the Stefan-Boltzmann limit, the behavior of thermal effective quark masses, and can reach possible regimes with causality violation Pasqualotto et al. (2023).
@@ II The NJL model and regularization
where the quark fields with two flavors $\psi=(u,d)^{T}$ are charge-conjugate spinors, with $\psi^{C}=C\bar{\psi}^{T}$, $\bar{\psi}^{C}=\psi^{T}C$ ($C=i\gamma^{2}\gamma^{0}$ is the charge conjugation matrix), and $\vec{\tau}=(\tau_{1},\tau_{2},\tau_{3})$ are the Pauli matrices. The current quark mass matrix and chemical potential matrices are defined as $\hat{m}=\text{diag}\left(m_{u},m_{d}\right)$, and $\hat{\mu}=\left(\mu_{ur},\mu_{ug},\mu_{ub},\mu_{dr},\mu_{dg},\mu_{db}\right)$, respectively. In the isospin-symmetric limit, we assume $m_{u}=m_{d}\equiv m_{c}$ and $\mu_{i}\equiv\mu$. Moreover, $\left(\epsilon_{c}^{3}\right)^{ab}=\left(\epsilon_{c}\right)^{3ab}$ and $\left(\epsilon_{f}\right)^{ij}$ are the antisymmetric matrices in color and flavor spaces, respectively. In this equation, one may see that the coupling of the quarks with the magnetic field $\tilde{\mathcal{A}}_{\mu}$ is implemented through the covariant derivative in terms of the rotated fields. In the presence of a non-vanishing superconducting gap $\Delta$, the photon acquires a finite mass Allen et al. (2015). However, as shown in Alford et al. (2000), there is a linear combination of the photon and a gluon field that remains massless. The associated rotated charge matrix $\tilde{\mathcal{Q}}$ is given by $\tilde{\mathcal{Q}}=\mathcal{Q}_{f}\otimes 1_{c}-1_{f}\otimes\left(\frac{\lambda_{8}}{2\sqrt{3}}\right)$, where $\mathcal{Q}_{f}=\text{diag}(2/3,-1/3)$ and $\lambda_{8}=\text{diag}(1,1,-2)/\sqrt{3}$ is the color quark matrix. Therefore, in the six-dimensional flavor-color representation, the rotated charges for the different quarks are $u_{r}=u_{g}=1/2,u_{b}=1,d_{r}=d_{g}=-1/2,d_{b}=0$. Considering a constant and static magnetic field in the 3-direction, $\tilde{A}_{\mu}=\delta_{\mu 2}x_{1}B$, we have a mixture of the electromagnetic field and color fields. The rotated unit charge $\tilde{e}=e\cos(\theta)$, where $\theta$ is the mixing angle, is estimated to be of the order of 1/20 Gorbar (2000). To simplify notation, we will use $e$ instead of $\tilde{e}$ in what follows. Lastly, $G_{S}$ and $G_{D}$ are the scalar-pseudoscalar and diquark coupling constants.
@@
$E^{\pm}_{p,1}=\sqrt{p_{z}+2eBn+M^{2}}\pm\mu\,$, (4)
$E^{\pm}_{p,\frac{1}{2}}=\sqrt{(\sqrt{p_{z}+eBn+M^{2}}\pm\mu)^{2}+\Delta^{2}},$ (5)
@@
In the definitions above, terms multiplying the Heaviside theta functions come from the $T=0$ limit of Eq. (2), while integrations over the dispersions are divergent and must be regularized. In the 2010s, it became common in the literature to use simple step functions, $\theta(x-\Lambda)$, or to introduce smooth cutoffs, often implemented through form factors, to regularize the $B$-dependent integrals as Fayazbakhsh and Sadooghi (2010, 2011); Fukushima and Warringa (2008b); Mandal and Jaikumar (2017)
@@
where $U\left(\sqrt{p_{z}^{2}+2|a|eBn}\right)$ is a smooth function that exhibits step-function-like behavior. Nevertheless, it has been shown that in both cases, this procedure introduces strong unphysical oscillations in the behavior of different quantities as functions of the magnetic field Avancini et al. (2019b). Although the use of smooth regulators improves the results compared to the simple step function, the unphysical oscillations are often incorrectly associated with van Alphen–de Haas (vA-dH) oscillations, even for parametrization sets where such oscillations should not be present Allen et al. (2015); Duarte et al. (2016, 2017). We will return to this point when discussing our numerical results.
In recent years, the importance of separating medium effects from the divergent integrals prior to regularization has been widely discussed. Regarding the magnetic field contributions, in this work, we employ the Magnetic Field Independent Regularization (MFIR), following the approach previously adopted by Allen et al. (2015); Duarte et al. (2016); Avancini et al. (2021). In the context of this work, the contributions $|a|=1$ and $|a|=1/2$ to the thermodynamic potential become (see Menezes et al. (2009); Allen et al. (2015) for explicit details of MFIR implementation):
@@
where the factor of 4 in front of the first integral in Eq. (10) differs from the factor of 2 in the first line of Eq. (8) due to an algebraic manipulation performed when applying MFIR. Specifically, the extra factor stems from the first term in $\Omega_{1}$ in the second line of Eq. (8). The remaining contribution of $\Omega_{1}$ reads:
@@
When using MFIR, we have divergent integrals of the type [^1] Please refer to Duarte et al. (2019) for the explicit form of the $\Delta$ and $M$ gap equations at zero magnetic field. The integral $I_{M,0}$ is the derivative of the first term in Eq. (10) with respect to $M$, while $I_{\Delta}$ and $I_{M,\frac{1}{2}}$ are the derivatives of the first line of Eq. (12) with respect to $\Delta$ and $M$, respectively. The extension to $eB\neq 0$ is straightforward.:
@@
for the $\Delta$ gap equation, for example. The integrand in Eq. (14) is obtained after integrating the left-hand side of Eq. (17) over $a$ from -$\infty$ to $\infty$. The details of the MSS implementation are thoroughly discussed in Refs. Farias et al. (2006, 2016); Duarte et al. (2019); Avancini et al. (2019a). Therefore, we restrict ourselves to presenting the final expressions here. For $\Delta$, we have:
@@
with the remaining definitions
@@
In the MSS framework, the thermodynamic potential becomes, after collecting the regularized integrals in both the first line of (10) and the first and second lines of (12):
@@
In this work, two separation schemes are employed: MFIR, which isolates purely magnetic contributions, and MSS, which isolates density contributions. Throughout the paper, MFIR refers solely to the magnetic separation, while the full disentanglement of medium effects is achieved through the combined implementation of MFIR and MSS. In this section, we presented the expressions for different regularization methods, along with references for readers interested in the detailed derivations. In the subsequent section, we present a comparative analysis of these methods through numerical results, followed by a discussion of their implications for physical quantities.

### v2

@@ Abstract
We analyze the impact of the medium separation scheme (MSS) on two-flavor color superconducting (2SC) dense quark matter under the influence of a constant external magnetic field. The effects of the proper treatment of the model divergences are examined through a comparison of different approaches, including the combined implementation of the magnetic field independent regularization (MFIR) and the MSS, as well as the standard use of smooth form factors. Our findings for the Nambu-Lasinio model emphasize the critical role of properly separating medium effects from vacuum contributions in the model. The combined MFIR-MSS scheme suppresses spurious unphysical oscillations, often misinterpreted in the literature as de Haas-van Alphen oscillations, and ensures the correct high-density behavior of the diquark condensate. Furthermore, within the MSS framework, the magnetization remains positive across the explored parameter space, in sharp contrast with the behavior obtained in the traditional approach.
@@ I Introduction
The possible existence of deconfined quark matter in the core of neutron stars Annala et al. (2020) has shed new light on the understanding of strongly interacting matter described by QCD at high densities. The expectation that quark matter becomes color superconducting at very high densities has been thoroughly reviewed in the literature Bailin and Love (1984). Perturbative calculations were later confirmed within a controlled weak-coupling framework in Ref. Son (1999), where the authors derived the parametric form of the superconducting gap. Nevertheless, these perturbative approaches are not applicable to the intermediate density regime, which is relevant for the phenomenology of compact objects. Therefore, alternative approaches must be developed to explore the properties of matter in this regime. From a theoretical perspective, much of our current understanding of the QCD phase structure largely stems from first-principles Monte Carlo simulations on the lattice Laermann and Philipsen (2003); Parotto (2023), although the well-known sign problem limits the applicability of this approach at finite chemical potential (see Nagata (2022) for a recent review). Consequently, our understanding of the behavior of cold and dense matter in the intermediate density regime relies heavily on the development of effective models. Chiral models such as the Nambu–Jona-Lasinio (NJL) Klevansky (1992); Buballa (2005) and the linear sigma model Ayala et al. (2015); Andersen and Nødtvedt (2025) aim to capture the essential features of strongly interacting matter by incorporating key symmetries and physical constraints of QCD, providing valuable insights into the equation of state and phase transitions relevant to compact stars and other dense systems.
@@
Color superconducting phases are particularly relevant in astrophysical contexts such as magnetars, where dense matter is subjected to strong magnetic fields. In these extreme environments, the standard electromagnetic Meissner effect may be absent for certain nonspherical superconducting phases, thereby allowing the magnetic flux to penetrate without significant energy cost Feng et al. (2010). Earlier investigations of magnetized color-superconducting quark matter have demonstrated that strong magnetic fields can significantly influence both pairing dynamics and the phase structure, modifying the behavior of the gap and the equation of state Abhishek and Mishra (2019). Within the two-flavor sector, rotated charges and anisotropic superconductivity have been investigated in Refs. Coppola et al. (2017); Cao and Zhuang (2015), emphasizing the necessity of incorporating electric and color neutrality constraints for more realistic descriptions. Furthermore, intense magnetic fields have been shown to induce oscillatory behavior in the gap parameters Fukushima and Warringa (2008a); Noronha and Shovkovy (2007), emphasizing the nontrivial interplay between magnetic fields and superconductivity. Beyond gap oscillations, more exotic phenomena may arise, such as the formation of inhomogeneous gluon condensates, which mitigate chromomagnetic instabilities in neutral two-flavor superconductors, and the emergence of rotated magnetic fields Ferrer and de la Incera (2006, 2007a); Yuan et al. (2024). These effects are crucial for accurately modeling compact stars, especially magnetars. In the three-flavor case, intense magnetic fields may induce new phases, such as the magnetic color-flavor-locked (MCFL) and Paramagnetic-Color-Flavor-Locked (PCFL) phases, wherein both the pairing structure and the spectrum of collective excitations are significantly modified Ferrer and de la Incera (2007b); Feng et al. (2011); Paulucci et al. (2011). The alignment of Cooper pair magnetic moments in these phases can lead to the formation of additional condensates, thereby enhancing the pairing energy in the presence of strong fields. From a theoretical perspective, renormalization-group analyses have elucidated the role of long-range magnetic interactions in color superconductivity at asymptotically high densities, leading to a non-BCS scaling of the superconducting gap, $\Delta\sim\mu\,g^{-5}\exp(-c/g)$ Son (1999); Hsu and Schwetz (2000), where $c$ is a constant. These insights are essential for the construction of consistent models of dense QCD matter under extreme magnetic conditions. Moreover, it has been proposed that strong background fields could induce an electromagnetic superconducting state of the vacuum through the condensation of charged vector $\rho$ mesons, suggesting an analogy between QCD and condensed matter systems Chernodub (2010, 2013). However, recent calculations based on extended NJL models Carlomagno et al. (2022a, b), which consistently include Schwinger phases and proper wave functions for $B\neq 0$, have found no evidence of such condensation in the range of magnetic fields considered, thus disfavoring the occurrence of this mechanism.
@@
In this work, we employ the NJL model to study diquark condensation in cold and dense quark matter under an external magnetic field. Given the nonrenormalizable nature of the NJL model, the regularization procedure must be handled with particular care. In this class of effective models, UV divergent momentum integrals must be handled, typically by introducing a sharp cutoff $\Lambda$, which is regarded as a model parameter fitted to physical observables. This cutoff defines an energy scale beyond which the model cannot be trusted. Standard procedures often mix vacuum and medium effects, introducing dependencies on chemical potential, temperature, and external fields. Given that the NJL model serves as an effective approximation of QCD, it is reasonable to impose that regularization affects only vacuum-dependent terms. This principle was first applied in studies of color superconductivity Farias et al. (2006) and has since been employed in different contexts to describe the QCD phase diagram. The separation, termed the medium separation scheme (MSS), contrasts with the traditional regularization scheme (TRS), which applies the cutoff to all divergent integrals that include physical quantities that depend on medium effects. One of the main motivations for this study is the evidence that the diquark condensate $\Delta$ increases with baryon chemical potential $\mu_{B}$. While lattice QCD simulations for $N_{c}=3$ are hindered by the sign problem Karsch (2002), studies for $N_{c}=2$ Kogut et al. (2001); Braguta et al. (2016), chiral perturbation theory (ChPT) Kogut et al. (2000) and the renormalization-group consistent treatment applied to the NJL model Gholami et al. (2025) show this increasing behavior of $\Delta$ with $\mu_{B}$. Using TRS, the diquark condensate starts to decrease at a value of chemical potential smaller than the energy scale $\Lambda$, and eventually vanishes. In contrast, employing MSS, a consistently increasing condensate is predicted, in agreement with other approaches.
@@
Over the recent past, significant attention has been devoted to the generation of magnetic fields in noncentral heavy-ion collisions, where the field strength can reach up to $10^{20}$ G, and which can significantly influence the QCD phases Bzdak and Skokov (2013); McLerran and Skokov (2014). The study of QCD under strong magnetic fields has expanded across different contexts, including magnetars, neutron star mergers, and early Universe physics. For comprehensive overviews of the phase structure of QCD in strong magnetic backgrounds, encompassing effective model approaches, lattice QCD results, and different physical contexts in which magnetic fields play a relevant role, see Refs. Andersen et al. (2016); Adhikari and others (2026); Endrodi (2025); Miransky and Shovkovy (2015); Bandyopadhyay and Farias (2021) and references therein. In this context, accurately modeling the behavior of quark matter under strong magnetic fields requires a careful choice of regularization schemes to avoid unphysical artifacts and ensure reliable predictions, especially when working with nonrenormalizable models. The choice of an appropriate regularization scheme is essential for accurately describing magnetized quark matter, as highlighted in previous studies Avancini et al. (2019b); Allen et al. (2015); Duarte et al. (2016, 2017). As an example, one may cite a recent work Endrődi and Markó (2019), where the authors fixed the running coupling of the two-flavor Polyakov extended NJL model (PNJL) by matching its predictions with LQCD results for the baryon spectrum in a strong magnetic background. This approach, which preserves the expected decrease of the pseudocritical temperature, $T_{pc}$, with the increase of the magnetic field for both chiral and deconfinement transitions, relies on Schwinger’s proper time formalism Schwinger (1951), where all integrals are regularized without separating vacuum and thermomagnetic contributions. However, while this method ensures consistency with LQCD predictions at moderate temperatures, it can lead to inconsistencies at higher temperatures or finite baryon chemical potentials, such as unphysical oscillations in magnetization and imaginary meson masses Costa et al. (2010a); Fayazbakhsh et al. (2012); Ruivo et al. (2010); Costa et al. (2010b). Different regularization schemes have been proposed to address these issues, with the most prominent being the magnetic field independent regularization (MFIR), proposed in Ebert et al. (1999); Ebert and Klimenko (2003) and applied in Menezes et al. (2009), which isolates vacuum divergences from medium contributions and removes unphysical oscillations Ebert and Klimenko (2003); Ebert et al. (1999). Additionally, variations in handling the thermal contribution, such as the standard proper time and the thermomagnetic regulated proper time schemes Endrődi and Markó (2019), affect the convergence to the Stefan-Boltzmann limit, the behavior of thermal effective quark masses, and can reach possible regimes with causality violation Pasqualotto et al. (2023).
@@ II The NJL model and regularization
where the quark fields with two flavors $\psi=(u,d)^{T}$ are charge-conjugate spinors, with $\psi^{C}=C\bar{\psi}^{T}$, $\bar{\psi}^{C}=\psi^{T}C$ ($C=i\gamma^{2}\gamma^{0}$ is the charge conjugation matrix), and $\vec{\tau}=(\tau_{1},\tau_{2},\tau_{3})$ are the Pauli matrices. The current quark mass matrix and chemical potential matrices are defined as $\hat{m}=\text{diag}\left(m_{u},m_{d}\right)$, and $\hat{\mu}=\left(\mu_{ur},\mu_{ug},\mu_{ub},\mu_{dr},\mu_{dg},\mu_{db}\right)$, respectively. In the isospin-symmetric limit, we assume $m_{u}=m_{d}\equiv m_{c}$ and $\mu_{i}\equiv\mu$. Moreover, $\left(\epsilon_{c}^{3}\right)^{ab}=\left(\epsilon_{c}\right)^{3ab}$ and $\left(\epsilon_{f}\right)^{ij}$ are the antisymmetric matrices in color and flavor spaces, respectively. In this equation, one may see that the coupling of the quarks with the magnetic field $\tilde{\mathcal{A}}_{\mu}$ is implemented through the covariant derivative in terms of the rotated fields. In the presence of a nonvanishing superconducting gap $\Delta$, the photon acquires a finite mass Allen et al. (2015). However, as shown in Alford et al. (2000), there is a linear combination of the photon and a gluon field that remains massless. The associated rotated charge matrix $\tilde{\mathcal{Q}}$ is given by $\tilde{\mathcal{Q}}=\mathcal{Q}_{f}\otimes 1_{c}-1_{f}\otimes\left(\frac{\lambda_{8}}{2\sqrt{3}}\right)$, where $\mathcal{Q}_{f}=\text{diag}(2/3,-1/3)$ and $\lambda_{8}=\text{diag}(1,1,-2)/\sqrt{3}$ is the color quark matrix. Therefore, in the six-dimensional flavor-color representation, the rotated charges for the different quarks are $u_{r}=u_{g}=1/2,u_{b}=1,d_{r}=d_{g}=-1/2,d_{b}=0$. Considering a constant and static magnetic field in the 3-direction, $\tilde{A}_{\mu}=\delta_{\mu 2}x_{1}B$, we have a mixture of the electromagnetic field and color fields. The rotated unit charge $\tilde{e}=e\cos(\theta)$, where $\theta$ is the mixing angle, is estimated to be of the order of 1/20 Gorbar (2000). To simplify notation, we will use $e$ instead of $\tilde{e}$ in what follows. Lastly, $G_{S}$ and $G_{D}$ are the scalar-pseudoscalar and diquark coupling constants.
@@
$E^{\pm}_{p,1}=\sqrt{p_{z}^{2}+2eBn+M^{2}}\pm\mu\,$, (4)
$E^{\pm}_{p,\frac{1}{2}}=\sqrt{(\sqrt{p_{z}^{2}+eBn+M^{2}}\pm\mu)^{2}+\Delta^{2}},$ (5)
@@
In the definitions above, terms multiplying the Heaviside theta functions come from the $T=0$ limit of Eq. (2), while integrations over the dispersions are divergent and must be regularized. In 2010, it became common in the literature to use simple step functions, $\theta(x-\Lambda)$, or to introduce smooth cutoffs, often implemented through form factors, to regularize the $B$-dependent integrals as Fayazbakhsh and Sadooghi (2010, 2011); Fukushima and Warringa (2008b); Mandal and Jaikumar (2013, 2016, 2017)
@@
where $U\left(\sqrt{p_{z}^{2}+2|a|eBn}\right)$ is a smooth function that exhibits step-function-like behavior. Nevertheless, it has been shown that in both cases, this procedure introduces strong unphysical oscillations in the behavior of different quantities as functions of the magnetic field Avancini et al. (2019b). Although the use of smooth regulators improves the results compared to the simple step function, the unphysical oscillations are often incorrectly associated with van Alphen–Haas (vA-dH) oscillations, even for parametrization sets where such oscillations should not be present Allen et al. (2015); Duarte et al. (2016, 2017). We will return to this point when discussing our numerical results.
In recent years, the importance of separating medium effects from the divergent integrals prior to regularization has been widely discussed. Regarding the magnetic field contributions, in this work, we employ the magnetic field independent regularization (MFIR), following the approach previously adopted by Allen et al. (2015); Duarte et al. (2016); Avancini et al. (2021). In the context of this work, the contributions $|a|=1$ and $|a|=1/2$ to the thermodynamic potential become (see Menezes et al. (2009); Allen et al. (2015) for explicit details of MFIR implementation),
@@
where the factor of 4 in front of the first integral in Eq. (10) differs from the factor of 2 in the first line of Eq. (8) due to an algebraic manipulation performed when applying the MFIR. Specifically, the extra factor stems from the first term in $\Omega_{1}$ in the second line of Eq. (8). The remaining contribution of $\Omega_{1}$ reads,
@@
When using MFIR, we have divergent integrals of the type [^1] Please refer to Duarte et al. (2019) for the explicit form of the $\Delta$ and $M$ gap equations at zero magnetic field. The integral $I_{M,0}$ is the derivative of the first term in Eq. (10) with respect to $M$, while $I_{\Delta}$ and $I_{M,\frac{1}{2}}$ are the derivatives of the first line of Eq. (12) with respect to $\Delta$ and $M$, respectively. The extension to $eB\neq 0$ is straightforward.,
@@
for the $\Delta$ gap equation, for example. The integrand in Eq. (14) is obtained after integrating the left-hand side of Eq. (17) over $a$ from -$\infty$ to $\infty$. The details of the MSS implementation are thoroughly discussed in Refs. Farias et al. (2006, 2016); Duarte et al. (2019); Avancini et al. (2019a). Therefore, we restrict ourselves to presenting the final expressions here. For $\Delta$, we have
@@
with the remaining definitions,
@@
In the MSS framework, the thermodynamic potential becomes, after collecting the regularized integrals in both the first line of (10) and the first and second lines of (12),
@@
In this work, two separation schemes are employed; MFIR, which isolates purely magnetic contributions, and MSS, which isolates density contributions. Throughout the paper, MFIR refers solely to the magnetic separation, while the full disentanglement of medium effects is achieved through the combined implementation of MFIR and MSS. In this section, we presented the expressions for different regularization methods, along with references for readers interested in the detailed derivations. In the subsequent section, we present a comparative analysis of these methods through numerical results, followed by a discussion of their implications for physical quantities.